# Asymptotically Stable Quaternion-valued Hopfield-structured Neural Networks with Periodic Projection-based Supervised Learning Rules

### v2

@@ 2 Background
Quaternion multiplication can be represented as real-valued matrix–vector multiplication over $\mathbb{R}^{4}$ by Eq. 2.1), where each quaternion defines a $4\times 4$ real matrix that acts on another quaternion interpreted as a 4-dimensional real vector. Due to the non-commutative nature of quaternion algebra, there exist two distinct but equivalent matrix representations: one for left multiplication and one for right multiplication. In this work, we adopt the left multiplication form consistently, where each quaternion induces a left-linear action on the quaternionic space. This operation is denoted by $\circ$ between two quaternions with the following form:
$\begin{split}\boldsymbol{q_{1}}\circ\boldsymbol{q_{2}}&=(s_{1}+x_{1}\boldsymbol{i}+y_{1}\boldsymbol{j}+z_{1}\boldsymbol{k})\circ(s_{2}+x_{2}\boldsymbol{i}+y_{2}\boldsymbol{j}+z_{2}\boldsymbol{k})\\ &=\left[\begin{matrix}s_{1}&-x_{1}&-y_{1}&-z_{1}\\ x_{1}&s_{1}&-z_{1}&y_{1}\\ y_{1}&z_{1}&s_{1}&-x_{1}\\ z_{1}&-y_{1}&x_{1}&s_{1}\\ \end{matrix}\right]\left[\begin{matrix}s_{2}\\ x_{2}\\ y_{2}\\ z_{2}\\ \end{matrix}\right]=\left[\begin{matrix}s_{2}&-x_{2}&-y_{2}&-z_{2}\\ x_{2}&s_{2}&z_{2}&-y_{2}\\ y_{2}&-z_{2}&s_{2}&x_{2}\\ z_{2}&y_{2}&-x_{2}&s_{2}\\ \end{matrix}\right]\left[\begin{matrix}s_{1}\\ x_{1}\\ y_{1}\\ z_{1}\\ \end{matrix}\right]\end{split}$ (2.1)
@@
$L(1)=\mathbb{I}_{4},L(i)=\left[\begin{matrix}0&1&0&0\\ -&0&0&0\\ 0&0&0&1\\ 0&0&-&0\end{matrix}\right],\ L(j)=\left[\begin{matrix}0&0&1&0\\ 0&0&0&-\\ -&0&0&0\\ 0&1&0&0\end{matrix}\right],\ L(k)=\left[\begin{matrix}0&0&0&1\\ 0&0&1&0\\ 0&-&0&0\\ -&0&0&0\end{matrix}\right].$ (2.2)
@@
And we also define the conjugate of a quaternion by $\boldsymbol{q}^{\dagger}=s-x\boldsymbol{i}-y\boldsymbol{j}-z\boldsymbol{k}=[s,-\boldsymbol{v}]$, the modulus or length of a quaternion by $|\boldsymbol{q}|=\sqrt{\boldsymbol{q}\circ\boldsymbol{q^{\dagger}}}=(s_{\boldsymbol{q}}^{2}+x_{\boldsymbol{q}}^{2}+y_{\boldsymbol{q}}^{2}+z_{\boldsymbol{q}}^{2})^{1/2}$, the inverse of a quaternion by $\boldsymbol{q}\circ\boldsymbol{q}^{-1}=1$, where $\boldsymbol{q}^{-1}=\boldsymbol{q}^{\dagger}/|\boldsymbol{q}|$. It is quite important to mention the geometric meaning and advantages of quaternion, which is just one of our motivations. There is a compact and intuitive form to represent 3-dimensional rotation by a quaternion.
@@
Rotating an arbitrary quaternion $\boldsymbol{q}=s_{\boldsymbol{q}}+\boldsymbol{v}_{\boldsymbol{q}}$ by a quaternion number parameterized by $\boldsymbol{\mu}=|\mu|(cos\beta+\boldsymbol{v}_{\boldsymbol{\mu}}sin\beta)$ is equivalent to computing the following:
@@
Definition 2.2 (Quaternion (left GHR) derivative Xu et al. (2015)).
Let $\boldsymbol{q}\in\mathbb{H}$ and $\boldsymbol{f}:D\rightarrow\mathbb{H},\ D\subseteq\mathbb{H}$, then the left GHR derivatives, with respect to $\boldsymbol{q^{\mu}}$ and $\boldsymbol{q^{\boldsymbol{\mu}}}$$^{*}$ ($\boldsymbol{\mu}\neq 0,\boldsymbol{\mu}\in\mathbb{H}$) of a well-defined function $\boldsymbol{f}$ are defined as:
@@ 3 Methodology
The application of GHR calculus.
Strict gradient descent is effective for the learning process, but the quaternion structure of the weights matrix $W$ could not be preserved, where we should have every $4\times 4$ block to be negative symmetric, except for the elements on the leading diagonal, like we have seen in the matrix expression of quaternionic multiplication by Eq. (2.1). GHR gradient descent over $\mathbb{H}$ can be solved from the following system. The first equation is the derivative of the loss function $E(\boldsymbol{q})=|\boldsymbol{q}-\boldsymbol{q}_{d}|^{2}$ through Def. 2.2, where $\boldsymbol{q}$ is the stable state of the evolution equation and $\mathcal{I}=\{{\boldsymbol{i},\boldsymbol{j},\boldsymbol{k}\}}$. The second equation is by taking the differential of the sensitivity Eq. (3.8). For $\varrho\in\mathcal{J}=\{1,\boldsymbol{i},\boldsymbol{j},\boldsymbol{k}\}$, the following differential could be calculated by:
@@ 6 Conclusions
The uniqueness of the system can be proved by inequality skills; the demand is converted to constraints of weights. Through reduction to absurdity, suppose there exist two different critical points $\boldsymbol{v}^{*},\ \boldsymbol{u}^{*}$. Consider $\boldsymbol{L_{1}}$ norm of vuv-u to be:
$||\boldsymbol{u},\boldsymbol{v}||_{1}$ $=\sum_{i=1}^{n}|u_{i}-v_{i}|=\sum_{i=1}^{n}\left|\frac{\mu}{\gamma}\sum_{j=1}^{n}w_{ij}[\varphi_{j}(u_{j})-\varphi_{j}(v_{j})]\right|\leq\sum_{i=1}^{n}\frac{\mu}{\gamma}\sum_{j=1}^{n}|w_{ij}|\cdot L_{i}\cdot|u_{j}-v_{j}|$ (A.4)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\mu}{\gamma}L_{i}\cdot|w_{ij}|\cdot|u_{j}-v_{j}|=\sum_{j=1}^{n}\left(\sum_{i=1}^{n}\frac{\mu}{\gamma}L_{i}|w_{ij}|\right)\cdot|u_{j}-v_{j}|<\sum_{j=1}^{n}|u_{i}-v_{i}|$
@@
For simplification, when we use the inequality A.5 for deeper deduction, we will suppose the coefficients in the front of weights to be a unit constant. And for the practical network operation, this condition will be satisfied during the procedure of weights normalization, where we set the norm of the weights matrix to be a constant and the significant information will not be lost.
@@
$\begin{split}\frac{\text{d}V(t)}{\text{d}t}&=\frac{1}{\gamma}\sum_{j=1}^{n}x_{j}\dot{x}_{j}\\ &=\frac{1}{\gamma}\sum_{j=1}^{n}x_{j}\left[-\gamma x_{j}+\mu\sum_{i=1}^{n}w_{ji}\varphi_{i}(x_{i})\right]\\[2.84544pt] &\leq\sum_{j=1}^{n}\left(-{x_{j}}^{2}+\frac{\mu}{\gamma}\sum_{i=1}^{n}|w_{ji}|L_{i}x_{i}x_{j}\right)\\[2.84544pt] &\leq\sum_{j=1}^{n}\left[-{x_{j}}^{2}+\frac{\mu}{2\gamma}\sum_{i=1}^{n}|w_{ji}|L_{i}({x_{i}}^{2}+{x_{j}}^{2})\right]\\[2.84544pt] &=\sum_{j=1}^{n}\left[-1+\frac{\mu}{2\gamma}\sum_{i=1}^{n}(|w_{ji}|+|w_{ij}|)\right]{x_{j}}^{2}\\[4.26773pt] &\leq 0\end{split}$ (B.3)
@@ Appendix D Proof of smoothness: Curvature bound for the smoothness
$\boldsymbol{z}_{i}:=\sum_{j}W_{ij}\phi(\boldsymbol{q}_{j})+\boldsymbol{b}_{j}$
@@
Now we will estimate the norm of different parts in D.3
@@
Combine D.6 and D.5 with D.3, we have a rough estimation of the 2nd derivative of $\boldsymbol{q}(t)$:
@@ Appendix F Baselines
Appendix G Absolute error to the target

### v3

# Asymptotically Stable Quaternionic Hopfield Structured Neural Network with Supervised Projection-based Manifold Learning

@@ 2 Background
Quaternion multiplication can be represented as real-valued matrix–vector multiplication over $\mathbb{R}^{4}$ by Eq. (2.2), where each quaternion defines a $4\times 4$ real matrix that acts on another quaternion interpreted as a 4-dimensional real vector. Due to the non-commutative nature of quaternion algebra, there exist two distinct but equivalent matrix representations: one for left multiplication and one for right multiplication. In this work, we adopt the left multiplication form consistently, where each quaternion induces a left-linear action on the quaternionic space. This operation is denoted by $\circ$ between two quaternions with the following form:
$\boldsymbol{q_{1}}\circ\boldsymbol{q_{2}}$ $=(s_{1}+x_{1}\boldsymbol{i}+y_{1}\boldsymbol{j}+z_{1}\boldsymbol{k})\circ(s_{2}+x_{2}\boldsymbol{i}+y_{2}\boldsymbol{j}+z_{2}\boldsymbol{k})$ (2.1)
$=\left[\begin{matrix}s_{1}&-x_{1}&-y_{1}&-z_{1}\\ x_{1}&s_{1}&-z_{1}&y_{1}\\ y_{1}&z_{1}&s_{1}&-x_{1}\\ z_{1}&-y_{1}&x_{1}&s_{1}\\ \end{matrix}\right]\left[\begin{matrix}s_{2}\\ x_{2}\\ y_{2}\\ z_{2}\\ \end{matrix}\right]=\left[\begin{matrix}s_{2}&-x_{2}&-y_{2}&-z_{2}\\ x_{2}&s_{2}&z_{2}&-y_{2}\\ y_{2}&-z_{2}&s_{2}&x_{2}\\ z_{2}&y_{2}&-x_{2}&s_{2}\\ \end{matrix}\right]\left[\begin{matrix}s_{1}\\ x_{1}\\ y_{1}\\ z_{1}\\ \end{matrix}\right]$ (2.2)
@@
$L(1)=\mathbb{I}_{4}\ \ L(\boldsymbol{i})=\left[\begin{matrix}0&-1&0&0\\ 1&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\end{matrix}\right]\ L(\boldsymbol{j})=\left[\begin{matrix}0&0&-1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&-1&0&0\end{matrix}\right]\ L(\boldsymbol{k})=\left[\begin{matrix}0&0&0&-1\\ 0&0&-1&0\\ 0&1&0&0\\ 1&0&0&0\end{matrix}\right]$ (2.3)
@@
And we also define the conjugate of a quaternion by $\boldsymbol{q}^{\dagger}=s-x\boldsymbol{i}-y\boldsymbol{j}-z\boldsymbol{k}=[s,-\boldsymbol{v}]$, the modulus or length of a quaternion by $|\boldsymbol{q}|=\sqrt{\boldsymbol{q}\circ\boldsymbol{q^{\dagger}}}=(s_{\boldsymbol{q}}^{2}+x_{\boldsymbol{q}}^{2}+y_{\boldsymbol{q}}^{2}+z_{\boldsymbol{q}}^{2})^{1/2}$, the inverse of a quaternion by $\boldsymbol{q}\circ\boldsymbol{q}^{-1}=1$, where $\boldsymbol{q}^{-1}=\boldsymbol{q}^{\dagger}/|\boldsymbol{q}|$. It is quite important to mention the geometric meaning and advantages of quaternion, which is just one of our motivations. There is a compact and intuitive form to represent 3-dimensional rotation by a quaternion. App. H provides more details on quaternion.
@@
Rotating an arbitrary quaternion $\boldsymbol{q}=s_{\boldsymbol{q}}+\boldsymbol{v}_{\boldsymbol{q}}$ by a quaternion number parameterized by $\boldsymbol{\mu}=|\mu|(cos\beta+\boldsymbol{v}_{\boldsymbol{\mu}}sin\beta)$ is equivalent to computing the following transformation:
@@
[…]
@@ 3 Methodology
[…]
@@ 6 Conclusions
$\leq\sum_{i=1}^{n}\frac{\mu^{2}}{\gamma^{2}}\Bigg[\left(\sum_{j=1}^{n}|w_{ij}|M\right)+|I_{i}|\Bigg]^{2}$ (A.4)
$:=\rho^{2}$ (A.5)
@@
The uniqueness of the system can be proved by inequality skills; the demand is converted to constraints of weights. Through reduction to absurdity, suppose there exist two different critical points $\boldsymbol{v}^{*},\ \boldsymbol{u}^{*}$. Consider $L_{1}$ norm of vuv-u to be:
$\|\boldsymbol{u},\boldsymbol{v}\|_{1}$ $=\sum_{i=1}^{n}|u_{i}-v_{i}|=\sum_{i=1}^{n}\left|\frac{\mu}{\gamma}\sum_{j=1}^{n}w_{ij}[\varphi_{j}(u_{j})-\varphi_{j}(v_{j})]\right|\leq\sum_{i,j=1}^{n}\frac{\mu}{\gamma}|w_{ij}|L_{i}|u_{j}-v_{j}|$ (A.6)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\mu}{\gamma}L_{i}|w_{ij}||u_{j}-v_{j}|=\sum_{j=1}^{n}\left(\sum_{i=1}^{n}\frac{\mu}{\gamma}L_{i}|w_{ij}|\right)|u_{j}-v_{j}|<\sum_{j=1}^{n}|u_{i}-v_{i}|$ (A.7)
@@
For simplification, when we use the inequality A.8 for deeper deduction, we will suppose the coefficients in the front of weights to be a unit constant. And for the practical network operation, this condition will be satisfied during the procedure of weights normalization, where we set the norm of the weights matrix to be a constant and the significant information will not be lost.
@@
$\frac{\text{d}V(t)}{\text{d}t}$ $=\frac{1}{\gamma}\sum_{j=1}^{n}x_{j}\dot{x}_{j}$ (B.3)
$=\frac{1}{\gamma}\sum_{j=1}^{n}x_{j}\left[-\gamma x_{j}+\mu\sum_{i=1}^{n}w_{ji}\varphi_{i}(x_{i})\right]$ (B.4)
$\leq\sum_{j=1}^{n}\left(-{x_{j}}^{2}+\frac{\mu}{\gamma}\sum_{i=1}^{n}|w_{ji}|L_{i}x_{i}x_{j}\right)$ (B.5)
$\leq\sum_{j=1}^{n}\left[-{x_{j}}^{2}+\frac{\mu}{2\gamma}\sum_{i=1}^{n}|w_{ji}|L_{i}({x_{i}}^{2}+{x_{j}}^{2})\right]$ (B.6)
$=\sum_{j=1}^{n}\left[-1+\frac{\mu}{2\gamma}\sum_{i=1}^{n}(|w_{ji}|+|w_{ij}|)\right]{x_{j}}^{2}$ (B.7)
$\leq 0$ (B.8)
@@ Appendix D Proof of smoothness: Curvature bound for the smoothness
$\boldsymbol{z}_{i}:=\sum_{j}W_{ij}\phi(\boldsymbol{q}_{j})+\boldsymbol{b}_{j}$ (D.2)
@@
Now we will estimate the norm of different parts in D.4
@@
Combine D.7 and D.6 with D.4, we have a rough estimation of the 2nd derivative of $\boldsymbol{q}(t)$:
@@ Appendix F Baselines
Appendix G Absolute error to the target with exponential decay
@@
Appendix H Quaternion, Lie group, Lie algebra, and matrix representation
Throughout the paper, it is noticeable that the matix representation of quaternion and its operation takes the central rule of algebraic embedding to the neural network. Here we provide more details and discussion of the background for this approach.
Quaternion division ring and Lie group. [12]
Let $\mathbb{H}^{\times}=\mathbb{H}\setminus\{0\}$ denote the set of nonzero quaternions. Equipped with quaternion multiplication $\circ$, $\mathbb{H}^{\times}$ is a (real) Lie group. Indeed, identifying $\mathbb{H}\cong\mathbb{R}^{4}$ as a real vector space shows that $\mathbb{H}^{\times}\cong\mathbb{R}^{4}\setminus\{0\}$ is an open submanifold, and the maps $(q,p)\mapsto q\circ p$ and $q\mapsto q^{-1}$ are smooth in real coordinates. The next paragraph gives a matrix representation of $\mathbb{H}^{\times}$, where we establish a Lie group homomorphism to a matrix Lie group $L_{q}$.
Matrix representation of quaternion multiplication. [33, 12]
Retrieve the notations in the main body, left multiplication mapping of $q\circ(\cdot)$ is denoted by $L(q):p\mapsto q\circ p$, as a linear mapping, the corresponding matrix representation exists and is denoted by $\mathcal{L}=\{L(q)\!:\!q\in\mathbb{H}\}\subset\mathbb{R}^{4\times 4}$. This set contains zero matrix $L(0)=O_{4\times 4}$, and it is closed under matrix multiplication. The invertible part $\mathcal{L}^{\times}=\mathcal{L}\setminus O_{4\times 4}=L(\mathbb{H}^{\times})\subset GL(4,\mathbb{R})$ is a matrix Lie group. Denote the vectorization by:
$\mathrm{Vec}(p):=(p_{0}\ p_{1}\ p_{2}\ p_{3})^{T}\in\mathbb{R}^{4},\quad\text{for }p=p_{0}+p_{1}\mathbf{i}+p_{2}\mathbf{j}+p_{3}\mathbf{k}.$ (H.1)
Then there exists a unique matrix $L(q)\in\mathbb{R}^{4\times 4}$ such that:
$\mathrm{Vec}(q\circ p)=L(q)\,\mathrm{Vec}(p),\quad\text{for all }p\in\mathbb{H}.$ (H.2)
A direct expansion of quaternion multiplication yields the explicit formula:
$L(q)\!=\!\!\left[\begin{matrix}q_{0}&-q_{1}&-q_{2}&-q_{3}\\ q_{1}&q_{0}&-q_{3}&q_{2}\\ q_{2}&q_{3}&q_{0}&-q_{1}\\ q_{3}&-q_{2}&q_{1}&q_{0}\end{matrix}\right]\!\!$ $=\!q_{0}\mathbb{I}_{4}\!+\!q_{1}\!\!\left[\begin{matrix}0&1&0&0\\ -&0&0&0\\ 0&0&0&1\\ 0&0&-&0\end{matrix}\right]\!\!+\!q_{2}\!\!\left[\begin{matrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{matrix}\right]\!\!+\!q_{3}\!\!\left[\begin{matrix}0&0&0&1\\ 0&0&1&0\\ 0&-&0&0\\ -&0&0&0\end{matrix}\right]$ (H.3)
$=q_{0}L(1)+q_{1}L(\boldsymbol{i})+q_{2}L(\boldsymbol{j})+q_{3}L(\boldsymbol{k})$ (H.4)
And it could be linearly decomposed by Eq. (2.3). The assignment $q\mapsto L(q)$ is a multiplicative representation in the sense that $L(q_{1}\circ q_{2})=L(q_{1})L(q_{2}),\ L(1)=\mathbb{I}_{4}$, for $q_{1},\ q_{2}\in\mathbb{H}^{\times}$ and hence it stablishes a Lie group homomorphism $\mathbb{H}^{\times}\to GL(4,\mathbb{R})$. Next we provide more properties on $L(q)$ and where the structure of Eq. (H.4) comes from.
[…]
Conjugation corresponds to transpose in this representation is $L(\overline{q})=L(q)^{T}$. Furthermore, $L(q)^{T}L(q)=L(\overline{q})L(q)=L(\overline{q}q)=L(\lVert q\rVert^{2})=\lVert q\rVert^{2}\mathbb{I}_{4}$. In particular, if $\lVert q\rVert=1$, then $L(q)\in SO(4)$, i.e., $L(q)$ is an orthogonal matrix and all $L(q)$ form a subgroup. For general $q\neq 0$, the matrix $L(q)$ is a scaling of an orthogonal transformation in $\mathbb{R}^{4}$, i.e.
$L(S^{3})\leq SO(4),\ \mathcal{L}^{\times}\cong L(S^{3})\times\mathbb{R}^{+}.$ (H.5)
where $S^{3}\!=\!\{q:\|q\|=1,q\in\mathbb{H}\}$ is the 3-dimensional spherical surface embedded in $\mathbb{R}^{4}$ consisting of unit quaternions according to the vectorization of $q$.
Skew-symmetric representation of imaginary quaternion and cross product. [33, 11]
Write $q=(q_{0},\mathbf{q})$ with $\mathbf{q}=(q_{1},q_{2},q_{3})^{T}\in\mathbb{R}^{3}$ as the vector component of quaternion. Define:
$\mathbf{q}]_{\times}=\left[\begin{matrix}0&-q_{3}&q_{2}\\ q_{3}&0&-q_{1}\\ -q_{2}&q_{1}&0\end{matrix}\right].$ (H.6)
Then $L(q)$ admits the decomposition in Eq. (H.4), we have the block splitting of $L(q)$:
$L(q)=\left[\begin{matrix}q_{0}&-\mathbf{q}^{T}\\ \mathbf{q}&q_{0}\mathbb{I}_{3}+[\mathbf{q}]_{\times}\end{matrix}\right].$ (H.7)
The mapping $\mathbf{q}\mapsto[\mathbf{q}]_{\times}$ is the isomorphism of $\mathbb{R}^{3}$, and it encodes the cross product via $\mathbf{q}]_{\times}\boldsymbol{x}=\mathbf{q}\times\boldsymbol{x}$ for all $x\in\mathbb{R}^{3}$. This explains why skew-symmetric structure naturally appears when representing quaternion multiplication by real matrices. Furthermore, matrix multiplication is compatible with the quaternion operation in matrix form, which is significant in computer programming.
Unit quaternion and rigid body rotation. [12, 11]
The set of unit quaternions $S^{3}$ is also a Lie group under quaternion multiplication. Consider the mapping $\pi:S^{3}\to SO(3)$ defined by group action on pure imaginary quaternions $\boldsymbol{x}\in\mathrm{Im}\,\mathbb{H}$ and $\boldsymbol{q}=\mathrm{cos}(\theta)+\mathrm{sin}(\theta)\hat{\boldsymbol{n}}\in S^{3}$, define the rotation transformation:
$\pi(\boldsymbol{q}):\boldsymbol{x}\mapsto\boldsymbol{q}\circ\boldsymbol{x}\circ\boldsymbol{q}^{-1}=R_{\hat{\boldsymbol{n}}}(\theta)\in SO(3).$ (H.8)
Notice that $\pi(\boldsymbol{q})=\pi(\boldsymbol{-q})$, and this mapping is a surjective Lie group homomorphism with kernel $\{\pm 1\}$, hence form a Lie group double covering of $SO(3)$. Thereby we have the relation:
$SO(3)\cong S^{3}/\{\pm 1\}\cong\mathbb{RP}^{3},$ (H.9)
by The First Isomorphism Theorem for Groups, which formalizes the fact that $\boldsymbol{q}$ and $\boldsymbol{q}$ represent the same 3D rotation. And $\mathbb{RP}^{3}$ is the real projective space of 3-dimension, defined by quotient space $(\mathbb{R}^{4}\setminus\{0\})/\sim$ under the equivalence relation $x\sim\lambda x$. Furthermore, considering the isomorphism of $S^{3}\cong SU(2)$, as observed from decomposition Eq. (H.4). We have the following relations on quaternion groups and rotation $SO(3)$:
$\quad S^{3}\cong SU(2)$ (H.10)
$\mathbb{H}^{\times}\cong SU(2)\times\mathbb{R}^{+}$ (H.11)
$SO(3$ $)\cong SU(2)/\{\pm\mathbb{I}\}\cong\mathbb{RP}^{3}$ (H.12)
These explain how we explain the dynamical evolution in the quaternion domain naturally by the rigid body rotation and directly corresponds to the joint pose of robotic manipulators.
Imaginary quaternion and Lie algebra. [12, 36]
The Lie algebra of $S^{3}$ can be identified with the tangent space at the identity element $e$ of Lie group $S^{3}$, which is the imaginary quaternions $(Im\mathbb{H},\circ)$. Thus, we have the following relation chain of algebraic isomorphism:
$\mathfrak{s}(3)\cong\mathfrak{su}(2)\cong\mathcal{T}_{e}S^{3}\cong\mathrm{Im}\,\mathbb{H}.$ (H.13)
Under this identification, the Lie bracket is realized by the quaternion commutator, and its role here is to connect the block representation $\mathbf{q}]_{\times}$ with the cross product. Concretely, for pure imaginary $\boldsymbol{u},\boldsymbol{v}\in\mathrm{Im}\,\mathbb{H}\cong\mathbb{R}^{3}$, define the communicator as the Lie bracket operation:
$\boldsymbol{u},\boldsymbol{v}]_{\mathbb{H}^{\times}}:=\boldsymbol{u}\!\circ\!\boldsymbol{v}\!-\!\boldsymbol{v}\!\circ\!\boldsymbol{u}=2\,\boldsymbol{u}\times\boldsymbol{v}\ \in\ \mathrm{Im}\,\mathbb{H}$ (H.14)
where the Lie bracket corresponds to the vector cross product (with the present normalization giving the factor $2$). Consider the Lie algebra of $SO(3)$ denoted by $\mathfrak{so}(3)$ to be the space of $3\times 3$ real skew-symmetric matrices $\{A\in\mathbb{R}^{3\times 3}:A^{T}\!=\!-A\}$, and the isomorphism $\mathbb{R}^{3}\cong\mathfrak{so}(3)$ is given by homomorphsim mapping $\mathbf{q}\mapsto[\mathbf{q}]_{\times}$. This is consistent with the matrix action used above: $\mathbf{q}]_{\times}\boldsymbol{x}=\mathbf{q}\times\boldsymbol{x}$. The Lie bracket operation of $\mathfrak{so}(3)$ is Eq. (H.15), given $X={[\boldsymbol{x}]}_{\times},\ Y={[\boldsymbol{y}]}_{\times}$, then
$X,Y]_{SO(3)}:=XY-YX=[\boldsymbol{x}\times\boldsymbol{y}]_{\times}\ \in\ \mathfrak{so}(3).$ (H.15)
Then we got the relation between these Lie algebra and cross product on $\mathbb{R}^{3}$:
$(\mathrm{Im}\,\mathbb{H},\ [\cdot,\cdot]_{\mathbb{H}^{\times}}/2)\cong(\mathbb{R}^{3},\times)\cong(\mathfrak{so}(3),\ [\cdot,\cdot]_{SO(3)})$ (H.16)
Isomorphisms $\mathbb{R}^{3}\to\mathfrak{so}(3)$ and $\mathfrak{s}(3)\to\mathbb{R}^{3}$ is defined by $\boldsymbol{x}\mapsto[\boldsymbol{x}]_{\times}$ and $\boldsymbol{q}\mapsto\mathrm{Vec}(\boldsymbol{q})$ respectively.
Remarks on the proposed model.
In this work, $L(q)$ is used as a structure-preserving matrix representation of quaternion multiplication, enabling quaternion-valued interactions to be written as real-valued block-matrix operations while retaining the exact algebraic coupling induced by $\mathbb{H}$. Importantly, the analysis and the learning dynamics in the main text do not require imposing the unit norm constraint $\|q\|=1$, where the rotational correspondense is decoded under projective meaning. And the representation $L(q)$ is well-defined for all $q\in\mathbb{H}$. By projection-based learning algorithm Alg. (1), we actually implement the process of manifold optimization on $\mathbb{H}^{\times}$. Instead of accurately calculating the tangent space of the manifold, our strategy could be understood as a first-order approximation of the Riemann gradient descent.